January 1, 2016, Lexo co bought fixed assets for $ 80,000. The expected life of the property, at the date of purchase, was five years and the residual value is estimated at $ 16,000. Lexo co uses a straight-line depreciation method for depreciation. At the beginning of the fourth year, the managers reassessed the value of the property and it now amounts to $ 8000. Based on this information, the depreciation (charges) at year-end 2019 should be:

a 12,800 $.
b 16,800 $.
c 33,600 $.
d 20,800 $.
Answer #1

Annual depreciation = (80000 - 16000)/5 = $12,800

Net book value at the beginning of 2019 = 80000 - 12800*3 = $41600

Depreciation (charges) at year-end 2019 = (41600-8000)*1/2 = $16,800

Option b. is correct answer.